OpenStudy (anonymous):

Find the area bounded by f(x)=sinx and the x-axis on [0, pi] using 4 partitions.

OpenStudy (anonymous):

which method? LHS, RHS, Midpoint Rule, Trapezoidal Rule ???

OpenStudy (anonymous):

RHS This is what I have so far: \[\Delta x = \frac{ \pi }{ 4 }\] \[A = \frac{ \pi }{ 4 } f(\frac{ \pi }{ 4 }) + \frac{ \pi }{ 4 } f(2(\frac{ \pi }{ 4 })) + \frac{ \pi }{ 4 } f(3(\frac{ \pi }{ 4 })) + \frac{ \pi }{ 4 } f(4(\frac{ \pi }{ 4 }))\] Am I on the right track?

OpenStudy (anonymous):

that looks good... :)

OpenStudy (anonymous):

simplified, you should have: \(\large A=\frac{\pi}{4}[sin(\frac{\pi}{4})+sin(\frac{2\pi}{4})+sin(\frac{3\pi}{4})+sin(\frac{4\pi}{4})] \)

OpenStudy (anonymous):

That matches what I got. Sweet! For an answer, I got ≈1.896.

OpenStudy (anonymous):

Thanks ByteMe for the guidance and assurance! I appreciate it!

OpenStudy (anonymous):

yw... :)